6. What are the units of k in the following rate law? M s2. B)- M2 C) M2 s M2 s. M3s Answer 7. Given the following rate law, how does of if the concentration of Yis doubled? the rate reaction change Rate k [x][Y]2 A) The rate of reaction will increase by a factor of 2. B) The rate of reaction will increase by a factor of 4. C) The rate of reaction will increase by a factor of 5. D) The rate of reaction will decrease by a factor of 2. E) The rate of reaction will remain unchanged. Answer:
Answer #1

6)
rate = k [X][Y]^2
put unit of all components except k,
M/s = k * M * M^2
M/s = k * M^3
k = M-2s-1
Answer: B

7)
order of Y is 2
so doubling Y will make rate 4 times
Answer: B

i am allowed to answer only 1 question at a time